# Cross-Lingual Learning vs. Low-Resource Fine-Tuning: A Case Study with Fact-Checking in Turkish

### 01 Mar 2024 v1

@@ 6.   Discussion
The main objective of this study is to test the possibility and the extent of making use of a large amount of fact-checking data and large language models that were heavily pre-trained in English for fact-checking in other languages with much less labeled data, and much smaller pre-training data for large language models. We focus on Turkish as a low-resource language for this task. Although focusing on a single familiar language allows us to curate a better fact-checking corpus, and perform more meaningful error analysis, our approach is applicable to many languages. Results are likely to differ based on typological similarity of the languages in question, as well other factors like geographical proximity and cultural similarity of the communities that speak the language.
@@
Our experiments demonstrate some small gains from the high-resource language in zero-shot and few-shot settings, where few-shot learning shows slight improvement over zero-shot. The results in Table 7 and Table 8 shows a small but consistent increase in F1-macro scores when a few examples are included. The benefit of more few-shot examples is unclear, however. The same is true for making use of machine translation from low-resource language to high-resource language. The test instances translated to English labeled by the models trained on English data clearly better than an uninformed system. However, transfer learning approaches compete with fine-tuning with the labeled low-resource language data. Even a small amount of training data provides better results than zero- or few-shot approaches.
@@
Another interesting outcome of our results is the success of small models that rely only on surface cues on the FCTR data. There are no obvious latent variables (e.g., authors, source websites) that can identify the veracity label of short claim texts. This means some relevant information is available on the surface features. However, the large language models surpass the simple ones on English with a large margin (see Table 3). This may indicate both the help of the linguistic and perhaps factual information brought by these models. [^10] A potential problem here is these models may have the full fact-checking report for the test instances, including the clearly stated verdict in their pre-training data. However, most probably the comparatively smaller Turkish data during pretraining is possibly a factor in low scores of Llama with fine-tuning with Turkish (Tables 4 and 5).
@@
In the majority of the experiments, only the claim statements were employed as input, since this is a more realistic scenario as individuals typically seek to assess the truthfulness of a claim before spending time gathering additional information. We also include evidence statements as input in some experiments, which show a clear benefit in providing additional information. However, evidence retrieval is also a challenging problem in fact-checking (which falls beyond the scope of this study). A further problem with providing evidence may be discouraging the model from leveraging its pre-trained knowledge while making decisions.
@@ 10.   Bibliographical References
  • Ebrahimi et al. (2018) Javid Ebrahimi, Anyi Rao, Daniel Lowd, and Dejing Dou. 2018. Hotflip: White-box adversarial examples for text classification. In Proceedings of the 56th Annual Meeting of the Association for Computational Linguistics (Volume 2: Short Papers), pages 31–36.
@@
  • Kazemi et al. (2022) Ashkan Kazemi, Zehua Li, Verónica Peréz-Rosas, Scott A Hale, and Rada Mihalcea. 2022. Matching tweets with applicable fact-checks across languages.
@@
  • Mangrulkar et al. (2022) Sourab Mangrulkar, Sylvain Gugger, Lysandre Debut, Younes Belkada, and Sayak Paul. 2022. Peft: State-of-the-art parameter-efficient fine-tuning methods. https://github.com/huggingface/peft.
  • Nakamura et al. (2020) Kai Nakamura, Sharon Levy, and William Yang Wang. 2020. Fakeddit: A new multimodal benchmark dataset for fine-grained fake news detection. In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 6149–6157.
@@
  • Schuster et al. (2019) Tal Schuster, Darsh Shah, Yun Jie Serene Yeo, Daniel Roberto Filizzola Ortiz, Enrico Santus, and Regina Barzilay. 2019. Towards debiasing fact verification models. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP), pages 3419–3425.
  • Shahi and Nandini (2020a) Gautam Kishore Shahi and Durgesh Nandini. 2020a. Fakecovid–a multilingual cross-domain fact check news dataset for covid-19. arXiv preprint arXiv:2006.11343.
  • Shahi and Nandini (2020b) Gautam Kishore Shahi and Durgesh Nandini. 2020b. Fakecovid–a multilingual cross-domain fact check news dataset for covid-19. arXiv preprint arXiv:2006.11343.
@@
  • Shu et al. (2020) Kai Shu, Deepak Mahudeswaran, Suhang Wang, Dongwon Lee, and Huan Liu. 2020. Fakenewsnet: A data repository with news content, social context, and spatiotemporal information for studying fake news on social media. Big data, 8(3):171–188.
@@
  • Zhang et al. (2019a) Chaowei Zhang, Ashish Gupta, Christian Kauten, Amit V Deokar, and Xiao Qin. 2019a. Detecting fake news for reducing misinformation risks using analytics approaches. European Journal of Operational Research, 279(3):1036–1052.
@@
  • Zhang et al. (2019b) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q Weinberger, and Yoav Artzi. 2019b. Bertscore: Evaluating text generation with bert. In International Conference on Learning Representations.
@@ Appendix A Topic Modeling
The topic distributions for each data split are given in Table 11 and Table 12 respectively. Even though we did not split the datasets according to the topic ratios, the most dominant and the least frequent topics were preserved in all data splits. For instance, in the FCTR dataset, The fifth topic is the most frequent topic in all subsets except FCTR500-val in which the given topic is not the most dominant topic by one margin. Additionally, the sixth topic is the least frequent topic in all splits.
Appendix B NELA Features

### 22 Mar 2024 v2

@@ 6.   Discussion
The main objective of this study is to test the possibility and the extent of making use of a large amount of fact-checking data and large language models that were heavily pretrained in English for fact-checking in other languages with much less labeled data, and much smaller pretraining data for large language models. We focus on Turkish as a low-resource language for this task. Although focusing on a single familiar language allows us to curate a better fact-checking corpus, and perform more meaningful error analysis, our approach is applicable to many languages. Results are likely to differ based on typological similarity of the languages in question, as well other factors like geographical proximity and cultural similarity of the communities that speak the language.
@@
Our experiments demonstrate some small gains from the high-resource language in zero-shot and few-shot settings, where few-shot learning shows slight improvement over zero-shot. The results in Table 7 and Table 8 shows a small but consistent increase in F1-macro scores when a few examples are included. The benefit of more few-shot examples is unclear, however. The same is true for making use of machine translation from low-resource language to high-resource language. The test instances translated to English labeled by the models trained on English data clearly better than an uninformed system. Even a small amount of training data provides better results than zero- or few-shot approaches.
@@
Another interesting outcome of our results is the success of small models that rely only on surface cues on the FCTR data. There are no obvious latent variables (e.g., authors, source websites) that can identify the veracity label of short claim texts. This means some relevant information is available on the surface features. However, the large language models surpass the simple ones on English with a large margin (see Table 3). This may indicate both the help of the linguistic and perhaps factual information brought by these models. [^10] A potential problem here is these models may have the full fact-checking report for the test instances, including the clearly stated verdict in their pretraining data. However, most probably the comparatively smaller Turkish data during pretraining is possibly a factor in low scores of LLaMA with fine-tuning with Turkish (Tables 4 and 5).
@@
In the majority of the experiments, only the claim statements were employed as input, since this is a more realistic scenario as individuals typically seek to assess the truthfulness of a claim before spending time gathering additional information. We also include evidence statements as input in some experiments, which show a clear benefit in providing additional information. However, evidence retrieval is also a challenging problem in fact-checking (which falls beyond the scope of this study). A further problem with providing evidence may be discouraging the model from leveraging its pretrained knowledge while making decisions.
@@ 10.   Bibliographical References
  • Kazemi et al. (2022) Ashkan Kazemi, Zehua Li, Verónica Pérez-Rosas, Scott A Hale, and Rada Mihalcea. 2022. Matching tweets with applicable fact-checks across languages. arXiv preprint arXiv:2202.07094.
@@
  • Nakamura et al. (2020) Kai Nakamura, Sharon Levy, and William Yang Wang. 2020. r/Fakeddit: A new multimodal benchmark dataset for fine-grained fake news detection. In Proceedings of the Twelfth Language Resources and Evaluation Conference, pages 6149–6157.
@@
  • Shahi and Nandini (2020) Gautam Kishore Shahi and Durgesh Nandini. 2020. FakeCovid–a multilingual cross-domain fact check news dataset for COVID-19. arXiv preprint arXiv:2006.11343.
@@
  • Zhang et al. (2019) Tianyi Zhang, Varsha Kishore, Felix Wu, Kilian Q Weinberger, and Yoav Artzi. 2019. BERTScore: Evaluating text generation with BERT. In International Conference on Learning Representations.
@@ Appendix A Topic Modeling
The topic distributions for each data split are given in Table 11 and Table 12 respectively. Even though we did not split the datasets according to the topic ratios, the most dominant and the least frequent topics were preserved in all data splits. For instance, in the FCTR dataset, The fifth topic is the most frequent topic in all subsets except FCTR500-val in which the given topic is not the most dominant topic by a small margin. Additionally, the sixth topic is the least frequent topic in all splits.
Appendix B NELA Features